Write the equation of the line through (2, 3) and perpendicular to y = -2x +
4 in slope-intercept form (y = mx + b).

Answer by anne(47)

ok

Perpendicular lines have opposite slopes
so our slope will be 1/2

and for the y intercept we have substitute the slope and two points.
(2,3)
so
y= mx+b
3= 1/2 (2) +b
3= 1+b
3-1= +1-1 +b
2=b


so our final equation is y= 1/2x +2


Did this help?
CHEERS